OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)

OpenStudy (lalaly):

first find the slope, do u know how?

OpenStudy (anonymous):

no

OpenStudy (lalaly):

ok the slope is equal to\[\large{\frac{\Delta y}{\Delta x}= \frac{y_2-y_1}{x_2-x_1}}\]

OpenStudy (lalaly):

\[=\frac{11-5}{-2-0}\]

OpenStudy (anonymous):

Lala, we can further simplify it to: \[ = -\frac{7}{2} \] :P

OpenStudy (lalaly):

oops 2

OpenStudy (anonymous):

No it's \[ =\frac{11-5}{-2-0} = -\frac{6}{2}\]

OpenStudy (lalaly):

\[\frac{-6}{2} = -3\]

OpenStudy (anonymous):

y – 5 = –3(x + 2) y – 5 = 3(x + 2) y – 11 = –3(x – 2) y – 11 = –3(x + 2) were the options though :/

OpenStudy (anonymous):

That's the hardest step lol

OpenStudy (lalaly):

yeah wer not done, i just showd u how to find the slope

OpenStudy (lalaly):

Do u want to learn this or do u just want the answer?

OpenStudy (anonymous):

Who wants to learn? :P here is the answer : \[ y – 11 = –3(x + 2) \]

OpenStudy (lalaly):

Lol Ok FOOOL :P

OpenStudy (anonymous):

haha :D

OpenStudy (anonymous):

thank you both *Bows down*

OpenStudy (anonymous):

;)